OpenStudy (anonymous):

What are the zeros of the polynomial function: f(x) = x3 + 9x2 + 20x ? 0, –5, –4 0, –5, 4 0, 5, –4 0, 5, 4

OpenStudy (anonymous):

Thanks again(: lol

OpenStudy (nurali):

Anytime.

OpenStudy (anonymous):

I'm sorry to point out that there was a mistake that you made Nurali. The original question had +9x^2. You did the math using -9x^2. The zeros are 0, -4, and -5.

OpenStudy (nurali):

f(x) = x^3 + 9x^2 + 20x x^3 + 9x^2 + 20x = 0 x(x^2 + 9x + 20) = 0 x(x + 4)(x + 5) = 0 x = 0 x + 4 = 0 x + 5 = 0 x = 0 x = -4 x = -5 the zeros are 0, -5, -4